# Extremal correlation coefficient for functional data

### v2

@@ 3 Bivariate regular variation in Banach spaces
This choice of norm works well with the extremal correlation coefficient defined in Section 4.
In order to derive the joint angular probability measure of $X$ and $Y$, we consider the polar coordinate transformation $T:{\mathbb{B}}_{0}^{2}\to\left([0,\infty)^{2}\setminus\{\mbox{\boldmath${0}$}\}\right)\times{\mathbb{S}}^{2}$, defined by
$T(x,y)=\left(\|x\|_{{\mathbb{B}}},\|y\|_{{\mathbb{B}}},\frac{x}{\|x\|_{{\mathbb{B}}}},\frac{y}{\|y\|_{{\mathbb{B}}}}\right)=:(r_{X},r_{Y},\theta_{X},\theta_{Y}),\ \ \ \ (x,y)\in{\mathbb{B}}_{0}^{2}.$ (7)
Using $T$, we obtain an equivalent formulation for a regularly varying random element in ${\mathbb{B}}^{2}$.
Proposition 3.2.
A bivariate random element $[X,Y]^{\top}$ in ${\mathbb{B}}^{2}$ is regularly varying with index $-\alpha$, $\alpha>0$, if and only if there exists an exponent measure $\nu$ in $M_{0}([0,\infty)^{2})$ and an angular probability measure $\Gamma$ in $M_{0}({\mathbb{S}}^{2})$ such that
$n{\rm pr}\left(\frac{(\|X\|_{{\mathbb{B}}},\|Y\|_{{\mathbb{B}}})}{b(n)}\in\cdot,\left(X/\|X\|_{{\mathbb{B}}},Y/\|Y\|_{{\mathbb{B}}}\right)\in\cdot\right)\stackrel{{\scriptstyle M_{0}}}{{\longrightarrow}}\nu\times\Gamma,\ \ n\to\infty,$ (8)
where $b(n)$ is the increasing sequence in (4). (We note that $\mu$ in (4) is a measure on ${\mathbb{B}}_{0}^{2}$, while $\nu$ in (8) is a measure on $[0,\infty)^{2}\setminus\{\mbox{\boldmath${0}$}\}$.)
Proof 3.3.
We only show that (4) implies (8) since showing the converse is similar. Take any $f\in{\mathcal{C}}_{0}([0,\infty)^{2}\times{\mathbb{S}}^{2})$. It then follows from the change of variables that
$\int_{[0,\infty)^{2}}\int_{{\mathbb{S}}^{2}}f(r_{X},r_{Y},\theta_{X},\theta_{Y})\ n{\rm pr}\left(\frac{(\|X\|_{{\mathbb{B}}},\|Y\|_{{\mathbb{B}}})}{b(n)}\in(dr_{X},dr_{Y}),\left(\frac{X}{\|X\|_{{\mathbb{B}}}},\frac{Y}{\|Y\|_{{\mathbb{B}}}}\right)\in(d\theta_{X},d\theta_{Y})\right)$
$=\int_{{\mathbb{B}}^{2}}f(T(x,y))\ n{\rm pr}\left(\frac{X}{b(n)}\in dx,\frac{Y}{b(n)}\in dy\right).$
Since $f\in{\mathcal{C}}_{0}([0,\infty)^{2}\times{\mathbb{S}}^{2})$, there exists a set $A$, bounded away from ${0}$ in $[0,\infty)^{2}\times{\mathbb{S}}^{2}$, such that $f(T(x,y))=0$ if $T(x,y)\notin A$. Then we have that $f(T(x,y))=0$ if $(x,y)\notin T^{-1}(A)$. Since $T^{-1}(A)$ is bounded away from ${0}$ in ${\mathbb{B}}^{2}$, we have that $f\circ T\in{\mathcal{C}}_{0}({\mathbb{B}}^{2})$. Then by the $M_{0}$ convergence in Definition 3.1, we have that
$\int_{{\mathbb{B}}^{2}}(f\circ T)(x,y)\ n{\rm pr}\left(\frac{X}{b(n)}\in dx,\frac{Y}{b(n)}\in dy\right)$
$\to\int_{{\mathbb{B}}^{2}}(f\circ T)(x,y)\mu(dx,dy)=\int_{T({\mathbb{B}}^{2})}f(r_{X},r_{Y},\theta_{X},\theta_{Y})\mu\circ T^{-1}(dr_{X},dr_{Y},d\theta_{X},d\theta_{Y}).$
To investigate the form of $\mu\circ T^{-1}$, take any $t>0$ and Borel set $S\subset{\mathbb{S}}^{2}$. It then follows from the homogeneity property of $\mu$ that
$\mu\circ T^{-1}([0,\infty)^{2}\setminus[0,t]^{2}\times S)$
$=\mu\left\{(x,y)\in{\mathbb{B}}^{2}_{0}:\|x\|_{{\mathbb{B}}}\vee\|y\|_{{\mathbb{B}}}>t,(x/\|x\|_{{\mathbb{B}}},y/\|y\|_{{\mathbb{B}}})\in S\right\}$
$=\mu\left\{(x,y)\in{\mathbb{B}}^{2}_{0}:\|x\|_{{\mathbb{B}}}\vee\|y\|_{{\mathbb{B}}}>t\right\}\times$
$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \frac{t^{-\alpha}\mu\left\{(x,y)\in{\mathbb{B}}^{2}_{0}:\|x\|_{{\mathbb{B}}}\vee\|y\|_{{\mathbb{B}}}>1,(x/\|x\|_{{\mathbb{B}}},y/\|y\|_{{\mathbb{B}}})\in S\right\}}{t^{-\alpha}\mu\left\{(x,y)\in{\mathbb{B}}^{2}_{0}:\|x\|_{{\mathbb{B}}}\vee\|y\|_{{\mathbb{B}}}>1\right\}}.$
It then follows from (6) that
$\mu\circ T^{-1}([0,\infty)^{2}\setminus[0,t]^{2}\times S)=\nu([0,\infty)^{2}\setminus[0,t]^{2})\Gamma(S),$
where
$\nu(A):=\mu\left\{(x,y)\in{\mathbb{B}}^{2}_{0}:(\|x\|_{{\mathbb{B}}},\|y\|_{{\mathbb{B}}})\in A\right\},\ \ \ \ A\subset[0,\infty)^{2}\setminus\{\mbox{\boldmath${0}$}\};$
$\Gamma(S):=\mu\left\{(x,y)\in{\mathbb{B}}^{2}_{0}:\|x\|_{{\mathbb{B}}}\vee\|y\|_{{\mathbb{B}}}>1,(x/\|x\|_{{\mathbb{B}}},y/\|y\|_{{\mathbb{B}}})\in S\right\},\ \ \ \ S\subset{\mathbb{S}}^{2}.$
Thus, $\mu\circ T^{-1}$ has the product form such that on $([0,\infty)^{2}\setminus\{\mbox{\boldmath${0}$}\})\times{\mathbb{S}}^{2}$
$\mu\circ T^{-1}=\nu\times\Gamma,$ (9)
which completes the proof.
Convergence (8) enables characterizing the joint extremal behavior of the bivariate regularly varying random vector $[X,Y]^{\top}$ in ${\mathbb{B}}^{2}$, using two measures: $\nu$ on $[0,\infty)^{2}\setminus\{\mbox{\boldmath${0}$}\}$ and $\Gamma$ on ${\mathbb{S}}^{2}$. The measure $\nu$ describes the joint extremal behavior of the scalars $\|X\|_{{\mathbb{B}}}$ and $\|Y\|_{{\mathbb{B}}}$. Specifically, if $\nu$ has its mass only on the axes, then $\|X\|_{{\mathbb{B}}}$ and $\|Y\|_{{\mathbb{B}}}$ are asymptotically independent, i.e., if one curve shows an extreme pattern, there is negligible probability of the other curve also showing an extreme pattern. If $\nu$ has mass only on the line $\{t(1,1),t>0\}$, then $\|X\|_{{\mathbb{B}}}$ and $\|Y\|_{{\mathbb{B}}}$ show asymptotic full dependence, i.e., extreme curves occur simultaneously in $X$ and $Y$.
The difference between $\nu$ in (8) and $\mu$ in (4) is that $\mu$ describes the joint behavior of extreme curves $X$ and $Y$ in ${\mathbb{B}}_{0}^{2}$, but $\nu$ describes extremal dependence between the sizes $\|X\|_{{\mathbb{B}}}$ and $\|Y\|_{{\mathbb{B}}}$ in $[0,\infty)^{2}\setminus\mbox{\boldmath${0}$}$. We remark that the measure $\nu$ on $[0,\infty)^{2}\setminus\mbox{\boldmath${0}$}$ is homogeneous like $\mu$ because for any $A\subset[0,\infty)^{2}\setminus\{\mbox{\boldmath${0}$}\}$ and $t>0$,
$\nu(tA)=\mu\left\{(x,y)\in{\mathbb{B}}^{2}_{0}:(\|x\|_{{\mathbb{B}}},\|y\|_{{\mathbb{B}}})\in tA\right\}=t^{-\alpha}\nu(A).$
The joint angular probability measure $\Gamma$ characterizes how the shapes of scaled $X$ and $Y$ are related in extremes. If the extreme curves are exactly proportional, i.e., $X=\lambda Y$, $\lambda\neq 0$, the scaled curves share the same extreme functional elements. This means that $\Gamma$ concentrates on the “line” $\{(\phi_{j},\phi_{j}),\ j\in{\mathcal{J}}\subset{\mathbb{N}}\}\subset{\mathbb{S}}^{2}$, where $\{\phi_{j},\ j\geq 1\}$ is a set of orthonormal elements in ${\mathbb{S}}$. If the shapes of two curves do not match in extremes, then $\Gamma$ concentrates on $\{(\phi_{j},\phi_{k})\}\subset{\mathbb{S}}^{2}$, where $j\in{\mathcal{J}}\subset{\mathbb{N}}$, $k\in{\mathcal{K}}\subset{\mathbb{N}}$, ${\mathcal{J}}\cap{\mathcal{K}}=\emptyset$. This situation corresponds to vanishing extremal covariance defined in Section 4.
@@
The marginal extreme behavior of $X$ can be obtained by integrating all possible values of $Y$ in (4), or $\|Y\|_{{\mathbb{B}}},Y/\|Y\|_{{\mathbb{B}}}$ in (8). Then $X$ has its marginal measure $\mu_{X}$, and equivalently $X\in RV(-\alpha,\Gamma_{X})$, where $\Gamma_{X}$ is the marginal angular measure of $X$. Similarly, $Y$ has its marginal $\mu_{Y}$, and equivalently $Y\in RV(-\alpha,\Gamma_{Y})$, where $\Gamma_{Y}$ is the marginal angular measure of $Y$.
@@ 4 Extremal correlation coefficient for functional data
In this section, we introduce the extremal correlation coefficient for functional data. It focuses on the extreme part of the joint distribution of regularly varying random elements $X$ and $Y$ in $L^{2}$ and measures the tendency of paired curves to exhibit similar extreme patterns.
To define the extremal correlation coefficient, we begin by introducing the concept of extremal covariance. Given a regularly varying bivariate random element $[X,Y]^{\top}$ in $L^{2}\times L^{2}$ with joint exponent measure $\mu$ in (4), we define the extremal covariance between $X$ and $Y$ by
@@
Recall that by (6), $\mu$ is a probability measure on the domain $\{(x,y)\in L^{2}\times L^{2}:\|x\|\vee\|y\|>1\}$, so $\mu I_{\|x\|\vee\|y\|>1}$ represents a probability distribution describing the joint extremal behavior of $X$ and $Y$. The extremal covariance is thus an extreme analog of the classic covariance in that $\sigma_{XY}$ measures how much two random curves vary together in extremes. We note that in order to define extremal covariance in (10), $[X,Y]^{\top}$ must be regularly varying with index $-\alpha$, where $\alpha>2$. The condition $\alpha>2$ is necessary because the definition of extremal covariance presumes the existence of the second moment, just as the usual covariance does. We elaborate on it at the end of this section, but note here that, as explained at the beginning of Section 5, regularly varying functions can be transformed to have the index $\alpha>2$.
Table 1: The range of $\sigma_{XY}$, extremal covariance between $X$ and $Y$, depending on the extremal dependence between $\|X\|$ and $\|Y\|$ and the level of similarity between $X/\|X\|$ and $Y/\|Y\|$ in extremes.
$\|X\|$ and $\|Y\|$ $\|X\|$ and $\|Y\|$
@@
$X/\|X\|$ and $Y/\|Y\|$
$\sigma_{XY}=0$     look similar $\sigma_{XY}>0$
@@
To examine the extremal covariance more closely, we recall the transformation $T$ defined in (7) and the relation $\mu\circ T^{-1}=\nu\times\Gamma$ in (9). It then follows from the change of variables that
$\sigma_{XY}=\int_{r_{X}\vee r_{Y}>1}r_{X}r_{Y}\ \nu(dr_{X},dr_{Y})\int_{{\mathbb{S}}^{2}}\left\langle\theta_{X},\theta_{Y}\right\rangle\Gamma(d\theta_{X},d\theta_{Y}).$ (11)
The extremal covariance of $X$ and $Y$ can be thus factorized into two components: the extremal dependence between $\|X\|$ and $\|Y\|$ (the first factor in (11)) and the level of similarity of the shapes between $X/\|X\|$ and $Y/\|Y\|$ (the second factor in (11)). These two components can be considered separately, each of them contains useful information, but there is value in combining them. If $\|X\|$ and $\|Y\|$ are asymptotically (in extremes) independent, then $\sigma_{XY}=0$ regardless of how the (scaled) extreme shapes of $X$ and $Y$ compare. The coefficient $\sigma_{XY}$ thus preserves the property that bivariate scalar observations are independent in extremes if they cannot take extremely large values at the same time. If “large” curves tend to occur simultaneously, i.e., $\|X\|$ and $\|Y\|$ are asymptotically dependent, then $\int_{r_{X}\vee r_{Y}>1}r_{X}r_{Y}\ \nu(dr_{X},dr_{Y})>0$. Consequently, there are three possible ranges for $\sigma_{XY}$, depending on the relative shape of extreme $X/\|X\|$ and $Y/\|Y\|$: 1) $\sigma_{XY}>0$ if the shapes look similar, 2) $\sigma_{XY}\approx 0$ if the shapes do not match, i.e., the curves are orthogonal, or 3) $\sigma_{XY}<0$ if they look opposite. These properties are summarized in Table 1. To provide a scale invariant measure, we define the extremal correlation coefficient.
@@
$\sigma_{X}=\left\{\int_{\|x\|\vee\|y\|>1}\|x\|^{2}\mu(dx,dy)\right\}^{1/2},\ \ \ \ \sigma_{Y}=\left\{\int_{\|x\|\vee\|y\|>1}\|y\|^{2}\mu(dx,dy)\right\}^{1/2}.$
The coefficient $\rho_{XY}$ has properties analogous to the classic correlation coefficient: 1) $-1\leq\rho_{XY}\leq 1$, 2) $\rho_{XY}$ measures the strength and direction of linear relationships between $X$ and $Y$ in extremes, 3) If $X$ and $Y$ are independent, then $\rho_{XY}=0$ since independence implies asymptotic independence between $\|X\|$ and $\|Y\|$.
To motivate our estimation approach, we first show that $\sigma_{XY}$ is a limit of the expected inner product of $X$ and $Y$ conditional on large values of $[X,Y]^{\top}$.
@@
Let $[X,Y]^{\top}$ be a regularly varying random element in $L^{2}\times L^{2}$. Then,
@@
$=\int_{L^{2}\times L^{2}}f(x,y)\frac{{\rm pr}({b(n)}^{-1}X\in dx,{b(n)}^{-1}Y\in dy)}{{\rm pr}(\|X\|\vee\|Y\|>{b(n)})}.$
Note that $f$ is bounded and vanishes on a neighborhood of ${0}$ in $L^{2}\times L^{2}$. Also, the discontinuity set of $f$ is the boundary of $\mathcal{A}_{1}=\{(x,y)\in L^{2}\times L^{2}:\|x\|\vee\|y\|\geq 1\}$, and it follows from Lemma A.1 that $\mu(\partial\mathcal{A}_{1})=0$. Therefore, by (4) and Lemma A.1 of Meinguet & Segers (2010), we get the claim.
@@
where $[X_{i},Y_{i}]^{\top},1\leq i\leq n$, are i.i.d. copies of $[X,Y]^{\top}$, $R_{i}:=\|X_{i}\|\vee\|Y_{i}\|$ and $R_{(k)}$ is the $k$th largest order statistic with the convention $R_{(1)}=\max\{R_{1},\ldots,R_{n}\}$. An estimator for $\rho_{XY}$ is then defined by
$\hat{\rho}_{n,k}=\frac{\sum_{i=1}^{n}\left\langle X_{i},Y_{i}\right\rangle}{\sqrt{\sum_{i=1}^{n}\|X_{i}\|^{2}}\sqrt{\sum_{i=1}^{n}\|Y_{i}\|^{2}}}\ I_{R_{i}\geq R_{(k)}}.$ (14)
These estimators take only the $k$ largest pairs of $[X_{i},Y_{i}]^{\top}$, $1\leq i\leq n$, based on their norm, i.e., $\|X_{i}\|\vee\|Y_{i}\|$, as inputs. This approach falls into so-called peaks-over-threshold framework in that it relies only on $k$ largest observations whose magnitude exceeds a certain threshold. Asymptotic properties in this framework are typically derived as $k$ goes to infinity with $n$, in such a way that $k/n\to 0$. We assume throughout the paper that this condition holds.
@@
The bivariate random element $[X,Y]^{\top}$ in $L^{2}\times L^{2}$ has mean zero and is regularly varying with index $-\alpha$, $\alpha>2$. The observations $[X_{1},Y_{1}]^{\top},[X_{2},Y_{2}]^{\top},\ldots$ are independent copies of $[X,Y]^{\top}$.
@@
From Theorem 4.4, the consistency of $\hat{\rho}_{n,k}$ for $\rho_{XY}$ follows from Slutsky’s theorem.
@@
We end this section with a discussion on the condition $\alpha>2$ in Assumption 4, which ensures the existence of the second moments, $E\|X\|^{2}$ and $E\|Y\|^{2}$. It can be dropped for the following alternative measure
$\gamma_{XY}:=\int_{{\mathbb{S}}^{2}}\left\langle\theta_{X},\theta_{Y}\right\rangle\Gamma(d\theta_{X},d\theta_{Y}),$
which is the angular density factor in (11). An estimator for $\gamma_{XY}$ can be defined by $\hat{\gamma}_{n,k}=\frac{1}{k}\sum_{i=1}^{n}\left\langle\frac{X_{i}}{\|X_{i}\|},\frac{Y_{i}}{\|Y_{i}\|}\right\rangle I_{R_{i}\geq R_{(k)}}$, and its consistency can be proven in almost the same manner as the proof of Corollary 4.2 of Clémençon et al. (2024). While $\gamma_{XY}$ can measure the similarity of curve shapes, it does not account for whether extreme curves $X$ and $Y$ occur simultaneously. Therefore, it can be argued that $\rho_{XY}$ is a more appropriate measure for evaluating concurrent risk despite requiring $\alpha>2$, as it assesses the similarity of shapes in paired extreme curves that occur simultaneously. To see this more closely, suppose that $X=Z_{1}\phi$ and $Y=Z_{2}\phi$, where $Z_{1}$, $Z_{2}$ are i.i.d. random variables satisfying ${\rm pr}(Z_{1}>z)=z^{-\alpha}$, $z>0$, and $\phi$ is any unit norm element of $L^{2}$. In this setting, two curves $X$ and $Y$ will appear identical in their extremes, but the extreme curves are unlikely to be observed simultaneously due to the independence of $Z_{1}$ and $Z_{2}$. Therefore, the assessment of concurrent risk should be zero, and the measure that gives zero is $\rho_{XY}=0$ (while $\gamma_{XY}=1)$. The identity $\rho_{XY}=0$ holds because the measure $\nu$ in (11) concentrates on the coordinate axes, so $\sigma_{XY}=\int_{r_{X}\vee r_{Y}>1}r_{X}r_{Y}\ \nu(dr_{X},dr_{Y})\int_{{\mathbb{S}}^{2}}\left\langle\theta_{X},\theta_{Y}\right\rangle\Gamma(d\theta_{X},d\theta_{Y})=0\times 1=0$, and the condition $\alpha>2$ is needed for the first integral to exist.

### v3

@@ 3 Bivariate regular variation in Banach spaces
This choice works well with the extremal correlation coefficient defined in Section 4.
Convergence (8) characterizes the joint extremal behavior of the bivariate regularly varying random vector $X,Y]^{\top}$ in ${\mathbb{B}}^{2}$ in terms of the tail index $\alpha$ and the joint angular measure $\Gamma$ on ${\mathbb{S}}_{{\mathbb{B}}^{2}}$. Here, $\alpha$ is the tail index of $\|X\|_{{\mathbb{B}}}\vee\|Y\|_{{\mathbb{B}}}$ and thus quantifies the frequency at which either $X$ or $Y$ exhibits an extreme size. The joint angular probability measure $\Gamma$ characterizes how the shapes of normalized $X$ and $Y$ relate in extremes. If the extreme curves are exactly proportional, i.e., $X=\lambda Y$, $\lambda\neq 0$, $\Gamma$ concentrates on pairs $(x,y)$ where the two components share the same functional shape up to scaling but with norms satisfying $\|x\|_{{\mathbb{B}}}\vee\|y\|_{{\mathbb{B}}}=1$. If the shapes of two curves differ substantially in extremes, $\Gamma$ places mass on pairs $(x,y)$ where the normalized curves exhibit distinct functional forms, such as orthogonality. This situation corresponds to vanishing extremal covariance defined in Section 4. The measure $\Gamma$ can also describe asymptotic independence between extremes of $X$ and $Y$, meaning that if one curve exhibits an extreme pattern, the probability that the other curve simultaneously exhibits an extreme pattern is negligible. In this case, $\Gamma$ concentrates on pairs $(x,y)$ where either $x=0$ or $y=0$, which also corresponds to vanishing extremal covariance.
@@
The marginal extremal behavior of $X$ can be obtained by integrating all possible values of $Y$ in (4). Then $X$ has its marginal measure $\mu_{X}$, and equivalently $X\in RV(-\alpha,\Gamma_{X})$, where $\Gamma_{X}$ is the marginal angular measure of $X$. Similarly, $Y$ has its marginal $\mu_{Y}$, and equivalently $Y\in RV(-\alpha,\Gamma_{Y})$, where $\Gamma_{Y}$ is the marginal angular measure of $Y$. We assume that the one-dimensional marginal distributions of $\mu$ are non-degenerate, i.e., $\mu_{X}$ and $\mu_{Y}$ are measures in $M_{0}({\mathbb{B}})$ satisfying analogs of (1).
@@ 4 Extremal correlation coefficient for functional data
In this section, we introduce the extremal correlation coefficient for functional data. It focuses on the extreme part of the joint distribution of regularly varying random vector $X,Y]^{\top}$ in $L^{2}\times L^{2}$ and measures the tendency of paired curves to exhibit similar extreme patterns.
To define the extremal correlation coefficient, we begin by introducing the concept of extremal covariance for functional data. Given a regularly varying bivariate random element $X,Y]^{\top}$ in $L^{2}\times L^{2}$ with joint exponent measure $\mu$ in (4), we define the extremal covariance between $X$ and $Y$ by
@@
Recall that by (7), $\mu$ is a probability measure on the domain $\{(x,y)\in L^{2}\times L^{2}:\|x\|\vee\|y\|>1\}$, so $\mu I_{\|x\|\vee\|y\|>1}$ represents a probability distribution describing the joint extremal behavior of $X$ and $Y$. The extremal covariance is thus an extreme analog of the classic covariance in that $\sigma_{XY}$ measures how much two random curves vary together in extremes. In order to define the extremal covariance, $X,Y]^{\top}$ must be regularly varying with index $\alpha$, where $\alpha>2$. The condition $\alpha>2$ is necessary because the definition of extremal covariance presumes the existence of the second moment, just as the usual covariance does. We elaborate on it at the end of this section, but note here that, as explained at the beginning of Section 5, regularly varying functions can be transformed to have the index $\alpha>2$.
To interpret the extremal covariance (10), we focus on the polar representation involving the joint angular measure $\Gamma$ in (9). Specifically, for any $(x,y)\neq(\mbox{${0}$},\mbox{${0}$})$, write $(x,y)=r(\theta_{X},\theta_{Y})$, where $r:=\|(x,y)\|$ and $(\theta_{X},\theta_{Y}):=(x,y)/\|(x,y)\|$. Then, $\mu$ can be decomposed as
$\mu(dx,dy)=\alpha r^{-\alpha-1}dr\Gamma(d\theta_{X},d\theta_{Y}).$
Using this, the extremal covariance $\sigma_{XY}$ can be expressed as
$\sigma_{XY}=\int_{r>1}r^{2}\alpha r^{-\alpha-1}dr\int_{{\mathbb{S}}_{L^{2}\times L^{2}}}\left\langle\theta_{X},\theta_{Y}\right\rangle\Gamma(d\theta_{X},d\theta_{Y})=\frac{\alpha}{\alpha-2}\int_{{\mathbb{S}}_{L^{2}\times L^{2}}}\left\langle\theta_{X},\theta_{Y}\right\rangle\Gamma(d\theta_{X},d\theta_{Y}).$ (11)
The extremal covariance of $X$ and $Y$ can thus be decomposed into the tail index factor $\alpha/(\alpha-2)$ and the angular dependence given by the integral.
Table 1: Ranges of $\sigma_{XY}$, the extremal covariance between $X$ and $Y$. Asymptotic independence refers to a joint distribution where extreme $X$ and $Y$ rarely occur simultaneously.
@@
$X/(\|X\|\vee\|Y\|)$ and $Y/(\|X\|\vee\|Y\|)$
$\sigma_{XY}\approx 0$    look similar $\sigma_{XY}>0$
@@
While the tail index factor represents the occurrence intensity of either extreme $\|X\|$ or $\|Y\|$, the angular dependence part captures the mutual relationship between the curves and how the shapes of the normalized extreme curves co-vary. More precisely, if the extremes of $X$ and $Y$ are asymptotically independent, meaning that extreme curves in $X$ and $Y$ rarely occur simultaneously, then $\Gamma$ places most mass near pairs where one component is close to the zero element, causing the integral to be close to zero, and thus $\sigma_{XY}$ will be near zero regardless of the shapes of the extreme curves. If $X$ and $Y$ are asymptotically dependent, meaning that extreme curves tend to occur simultaneously in $X$ and $Y$, then the value of $\sigma_{XY}$ depends on the shape similarity between the normalized curves $X/(\|X\|\vee\|Y\|)$ and $Y/(\|X\|\vee\|Y\|)$. Consequently, there are three possible ranges for $\sigma_{XY}$, depending on the relative shape of these normalized extremes: (i) $\sigma_{XY}>0$ when the shapes are similar, (ii) $\sigma_{XY}\approx 0$ when the shapes do not match and are approximately orthogonal, or (iii) $\sigma_{XY}<0$ when the shapes are opposite. These properties are summarized in Table 1.
To provide a scale-invariant measure of extremal dependence, we next define the extremal correlation coefficient.
@@
$\sigma_{X}=\left(\int_{\|x\|\vee\|y\|>1}\|x\|^{2}\mu(dx,dy)\right)^{1/2},\ \ \ \ \sigma_{Y}=\left(\int_{\|x\|\vee\|y\|>1}\|y\|^{2}\mu(dx,dy)\right)^{1/2}.$
The coefficient $\rho_{XY}$ has properties analogous to the classic correlation coefficient: (i) $1\leq\rho_{XY}\leq 1$, (ii) $\rho_{XY}$ measures the strength and direction of linear relationships between $X$ and $Y$ in extremes, (iii) if $X$ and $Y$ are independent, then $\rho_{XY}=0$ since independence implies asymptotic independence between $\|X\|$ and $\|Y\|$. We note that $\rho_{XY}$ may depend on $\alpha$ if $\Gamma$ depends on $\alpha$. An example is provided in Lemma 5.1, see also Example 2.2.8 in Kulik & Soulier (2020). If $\Gamma$ does not depend on $\alpha$, then $\rho_{XY}$ does not depend on $\alpha$.
To motivate our estimation approach, we first show that $\sigma_{XY}$ is a limit of the expected inner product of scaled $X$ and $Y$ conditional on large values of $X,Y]^{\top}$.
@@
Let $X,Y]^{\top}$ be a regularly varying random element in $L^{2}\times L^{2}$. Then,
@@
$=\frac{1}{n{\rm pr}(\|X\|\vee\|Y\|>{b(n)})}\int_{L^{2}\times L^{2}}f(x,y)\ n{\rm pr}({b(n)}^{-1}X\in dx,{b(n)}^{-1}Y\in dy).$
By (7), we have that $n{\rm pr}(\|X\|\vee\|Y\|>{b(n)})\to 1$. The function $f$ vanishes on a neighborhood of $(\mbox{${0}$},\mbox{${0}$})$ in $L^{2}\times L^{2}$. Also, the discontinuity set of $f$ is the boundary of $\mathcal{A}_{1}=\{(x,y)\in L^{2}\times L^{2}:\|x\|\vee\|y\|\geq 1\}$, and it follows from Lemma A.1 that $\mu(\partial\mathcal{A}_{1})=0$. Therefore, by (4) and an extension of Lemma A.1 of Meinguet & Segers (2010) (see Lemma A.6), we get the claim.
@@
where $X_{i},Y_{i}]^{\top},1\leq i\leq n$, are i.i.d. copies of $X,Y]^{\top}$, $R_{i}:=\|X_{i}\|\vee\|Y_{i}\|$ and $R_{(k)}$ is the $k$th largest order statistic with the convention $R_{(1)}=\max\{R_{1},\ldots,R_{n}\}$. An estimator for $\rho_{XY}$ is then defined by
$\hat{\rho}_{n,k}=\frac{\sum_{i=1}^{n}\left\langle X_{i},Y_{i}\right\rangle}{(\sum_{i=1}^{n}\|X_{i}\|^{2})^{1/2}(\sum_{i=1}^{n}\|Y_{i}\|^{2})^{1/2}}\ I_{R_{i}\geq R_{(k)}}.$ (14)
These estimators take only the $k$ largest pairs of $X_{i},Y_{i}]^{\top}$, $1\leq i\leq n$, based on their norm, i.e., $\|X_{i}\|\vee\|Y_{i}\|$, as inputs. This approach falls into so-called peaks-over-threshold framework in that it relies only on $k$ largest observations whose magnitude exceeds a certain threshold. Asymptotic properties in this framework are typically derived as $k$ goes to infinity with $n$, in such a way that $k/n\to 0$. We assume throughout the paper that this condition holds.
@@
The bivariate random element $X,Y]^{\top}$ in $L^{2}\times L^{2}$ has mean zero and is regularly varying with index $\alpha$, $\alpha>2$. The observations $X_{1},Y_{1}]^{\top},[X_{2},Y_{2}]^{\top},\ldots$ are independent copies of $X,Y]^{\top}$.
@@
The consistency of $\hat{\rho}_{n,k}$ for $\rho_{XY}$ follows from Theorem 4.4 and Slutsky’s theorem.
@@
We end this section with a discussion on the condition $\alpha>2$ in Assumption 4. This requirement ensures the existence of the second moments, $E\|X\|^{2}$ and $E\|Y\|^{2}$, allowing us to define $\sigma_{XY}$. If one wishes to avoid the condition on $\alpha$, the following alternative measure can be considered:
$\gamma_{XY}:=\int_{{\mathbb{S}}_{L^{2}\times L^{2}}}\left\langle\theta_{X},\theta_{Y}\right\rangle\Gamma(d\theta_{X},d\theta_{Y}),$
which corresponds to the angular density factor in (11). This measure depends only on the angular measure $\Gamma$, and is therefore defined for any $\alpha>0$. It can be regarded as an extension of the extremal dependence measure of Larsson & Resnick (2012) to paired functional samples, since their measure is also based on an angular measure. An estimator for $\gamma_{XY}$ can be defined as $\hat{\gamma}_{n,k}=\frac{1}{k}\sum_{i=1}^{n}\left\langle X_{i}/R_{i},Y_{i}/R_{i}\right\rangle I_{R_{i}\geq R_{(k)}}$, and its consistency can be proven in almost the same manner as the proof of Corollary 4.2 of Clémençon et al. (2024). The measure $\gamma_{XY}$ provides a convenient way to quantify the similarity in the shapes of extreme curves without requiring $\alpha>2$. However, $\gamma_{XY}$ does not account for the frequency of occurrence of extreme $X$ or $Y$ as $\sigma_{X,Y}$ does.